OpenStudy (anonymous):

The probability that a vehicle entering the Luray Caverns has Canadian license plates is 0.12; the probability that it is a camper is 0.28; and the probability that it is a camper with Canadian license plates is 0.09. What is the probability that (a) a camper entering the Luray Caverns has Canadian license plates? (b) a vehicle with Canadian license plates entering the Luray Caverns is a camper? (c) a vehicle entering the Luray Caverns does not have Canadian plates or is not a camper?

OpenStudy (anonymous):

So you have \[P(A)=P(\text{Canadian})=0.12\\ P(B)=P(\text{camper})=0.28\\ P(A\cap B)=P(\text{Canadian}\cap\text{camper})=0.09\] The first question wants to find the probability of a Canadian driver given that he is a camper: \[P(A|B)=\frac{P(A \cap B)}{P(B)}=\frac{0.09}{0.28}\] The second question asks for the probability of a camper given that he is Canadian: \[P(B|A)=\frac{P(B\cap A)}{P(A)}=\cdots\] The third wants the probability of "not a camper" OR "not Canadian," which translates to \[P(A^C\cup B^C)\] where \(A^C\) is the complement of \(A\). You can use deMorgan's laws to rewrite it: \[A^C\cup B^C=(A\cap B)^C\] And since this event and *its* complement (which is just \(A\cap B\)) are mutually exclusive, the probability of one occurring is given by the following relationships: \[P(A\cap B)=1-P\left((A\cap B)^C\right)\\ \color{red}{P\left((A\cap B)^C\right)}=1-P(A\cap B)\] Solve for the red part.

OpenStudy (anonymous):

Thank you so much! =)

OpenStudy (anonymous):

yw!